If x and y are real, solve the equation
jx/1+jy=3x+4j/x+3y
1
Expert's answer
2021-10-21T13:28:30-0400
"\dfrac{jx}{1+jy}=\dfrac{3x+4j}{x+3y}"

"j(x^2+xy)=3x-4y+j(xy+4)"

If "x" and "y" are real


"\begin{cases}
 3x-4y=0 \\
 x^2+xy=xy+4
\end{cases}"

"x=-2, y=-1.5"

"x=2, y=1.5"


"(-2, -1.5),\ (2,1.5)"